OpenStudy (anonymous):

Tickets for a play cost $3 for adults and $2 for children. 350 tickets were sold and $950 dollars was collected. If I wanted to find out how many tickets were sold what would my two equations be?

OpenStudy (colouringme):

Mathematical sentences (:

OpenStudy (anonymous):

I know that but what would the mathematical sentence look like. The first one would be 3a+2c=950d but what would the other look like?

OpenStudy (colouringme):

then a would represent what?

OpenStudy (colouringme):

you won't need d

OpenStudy (colouringme):

first you should know what does a and c represent.

OpenStudy (anonymous):

Let x be the no. of Adults who bought the ticket and y be no. of children who bought the ticket So We Have x+y=total no. of people=total no. of tickets sold=350 and 3x+2y=950 Njoy :)

OpenStudy (anonymous):

Basically x+y=350 and 3x+2y=950

OpenStudy (anonymous):

Thank you so much